A tortoise can run with a speed of 0.10 m/s, and a hare can run 20 times as fast. In a race, they both start at the same time, but the hare stops to rest for 5.0 minutes. The tortoise wins by a shell (20 cm).

a) How long does the race take?

b) What is the length of the race?
Answer #1

fine Velocity of tortoise, U = 0.10 m/s velocity of hare » Vn=20x0. = 2000-1 me 5 min = 5x60 see time = 2 m/s 2300 see - 2oom